The length of the top of a table is 7m greater than the width. The aread is 78m^2 Find the dimensions of the table. What is the width and length of the table?
Answer by stanbon(75887)

The length of the top of a table is 7m greater than the width.
Let width be "x"
Then length is "x+7"
-----------------------------
The area is 78m^2 Find the dimensions of the table.
Equation:
width*length = area
x(x+7) = 78
x^2+7x-78 = 0
---
(x-6)(x+13) = 0
Positive solution:
x = 6m (width)
---
x+7 = 13m (length)
=====================
Cheers,
Stan H.
=========

What is the width and length of the table?